Solve for x
y=w/x-2

Answer by chriswen(106)

I'm assuming you mean solve for x in terms of y and w.
y = w/x - 2
Add two on both sides
y + 2 = w / x
Multiply both sides by x
x ( y + 2 ) = w
Divide both sides by y + 2
x = w / (y + 2)